OpenStudy (facebook):

200 divide by 40 plus 4 times 3

OpenStudy (anonymous):

27

OpenStudy (facebook):

how

OpenStudy (anonymous):

do the problem

OpenStudy (anonymous):

on a calculator

OpenStudy (facebook):

kk

OpenStudy (facebook):

cant school work bu

OpenStudy (facebook):

just kidding

OpenStudy (anonymous):

its 27 200/40=5, 5+4=9, 9x3=27

OpenStudy (facebook):

but i do need help on it draw me a pic

OpenStudy (anonymous):

Equation: 200 / 40 + 4 x 3 50 + 4 x 3 50 + 12 Answer: 62

OpenStudy (anonymous):

howd you get 62?

OpenStudy (facebook):

what 62

OpenStudy (anonymous):

200/40=5 5+4=9 9x3=27

OpenStudy (anonymous):

@CodyKeller

OpenStudy (anonymous):

PEMDAS Parentheses Exponents Multiplication Division Addition Subtraction You need to do these in order.

OpenStudy (anonymous):

200 / 40 = 50 not 5

OpenStudy (anonymous):

thats 5

OpenStudy (anonymous):

Wow. Lol, sorry, ( feels stupid ) small mathematical error...

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

i was like ummmm, so it is 27 @facebook

OpenStudy (anonymous):

Lol. Sorry for the confusion.

OpenStudy (anonymous):

lol dont worry about it

OpenStudy (anonymous):

i dont even think @facebook is here..

OpenStudy (anonymous):

Wait, I just got 17 not 27

OpenStudy (anonymous):

you need to do the multiplication before the addition

OpenStudy (anonymous):

How? The rules of pemdas dont follow in this math problem

OpenStudy (anonymous):

They dont?

OpenStudy (anonymous):

No, well not the way i was tough .

OpenStudy (anonymous):

Well i dont think it matter because facebook put down 62